Check whether the value given in the brackets is a solution to the given equation or not:
4p – 3 = 13 (p = 0)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = 4p - 3
By substituting the value of p = 0
Then,
LHS = 4p - 3
= (4 × 0) - 3
=0 - 3
= -3
By comparing LHS and RHS
- 3≠13
LHS≠RHS
Hence, the value of p = 0 is not a solution to the given equation 4p - 3 = 13.

Thanks